Expand and simplify:
a. 6(x+2)+(x+2)
b. 5(2x-1)-3x(2x-1)

Answer by HyperBrain(694)

a. 6(x+2)+(x+2)
b. 5(2x-1)-3x(2x-1)
do you remember this property?
ax+bx=x(a+b)
let's apply this property
a.6(x+2)+(x+2)
6(x+2)+(x+2)=6(x+2)+1(x+2)
=(6+1)(x+2)
=7(x+2)
=7x+7*2=7x+14
b. 5(2x-1)-3x(2x-1)
=(5-3x)(2x-1)
=(-3x+5)(2x-1)
=-6x2 +13x -5
Power up,
HyperBrain!